# DS-Agent: Automated Data Science by Empowering Large Language Models with Case-Based Reasoning

### v3

@@ Abstract
In this work, we investigate the potential of large language models (LLMs) based agents to automate data science tasks, with the goal of comprehending task requirements, then building and training the best-fit machine learning models. Despite their widespread success, existing LLM agents are hindered by generating unreasonable experiment plans within this scenario. To this end, we present DS-Agent, a novel automatic framework that harnesses LLM agent and case-based reasoning (CBR). In the development stage, DS-Agent follows the CBR framework to structure an automatic iteration pipeline, which can flexibly capitalize on the expert knowledge from Kaggle, and facilitate consistent performance improvement through the feedback mechanism. Moreover, DS-Agent implements a low-resource deployment stage with a simplified CBR paradigm to adapt past successful solutions from the development stage for direct code generation, significantly reducing the demand on foundational capabilities of LLMs. Empirically, DS-Agent with GPT-4 achieves an unprecedented 100% success rate in the development stage, while attaining 36% improvement on average one pass rate across alternative LLMs in the deployment stage. In both stages, DS-Agent achieves the best rank in performance, costing \$1.60 and \$0.13 per run with GPT-4, respectively. Our code is open-sourced at https://github.com/guosyjlu/DS-Agent.
@@ 1 Introduction
Empirically, we demonstrate the superiority of DS-Agent across 30 data science tasks in both stages. During the development stage, DS-Agent with GPT-4 achieves an unprecedented 100% success rate across 12 tasks. For the deployment stage, DS-Agent with GPT-3.5 and GPT-4 achieves 85% and 99% one pass rate over 18 deployment tasks, while the best baseline attains a mere 56% and 60%. Remarkably, DS-Agent improves the one pass rate of an open-source LLM Mixtral-8x7b-Instruct from a mere 6% to 31%. In both stages, DS-Agent with GPT-4 and GPT-3.5 attains the highest and second-highest ranks in performance. Moreover, DS-Agent costs \$0.06 and \$1.60 per run for GPT-3.5 and GPT-4 with standard scenario, which reduce further to \$0.0045 and \$0.135 in low-resource scenarios, rendering DS-Agent highly appealing for real-world deployment.
@@ 3.1 Development Stage: Automatic Iteration Pipeline
We summarize the pseudo-code of the automatic pipeline in Algorithm 1. Overall, DS-Agent benefits from the CBR paradigm in two aspects. Firstly, CBR integrates the human insight case bank, which contains intensive expert knowledge of the data science, enabling DS-Agent to derive reasonable experiment plans. Moreover, CBR offers a flexible learning mechanism by retaining successful solution cases into the human insight case bank, thus eliminating the need for resource-intensive finetuning of LLMs via back-propagation. For instance, when encountering novel tasks that involve previously unseen data modalities, such as graph data, it can simply integrate the latest human insights into the case bank $\mathcal{C}$. This enables DS-Agent to adeptly solve data science tasks related to graph data by drawing on its enriched knowledge repository.
@@ 4.1 Experiment Setting
Evaluation Metric. We mainly evaluate the agent’s ability from three aspects: (1) Completion of building ML models. In the development stage, we employ the success rate, i.e., whether the agent can build an ML model within a fixed number of steps. In the deployment stage, the metric is the one-pass rate, indicating the agent’s ability to build an ML model with only a single trial. (2) Performance of built ML models. For both stages, we utilize mean rank and best rank as the evaluation metric to evaluate the agents’ capabilities for automated data science. (3) Resource cost. Since we mainly utilize closed-source LLMs in this work, we take the consumed money to assess resource costs.
@@ 4.2.1 Main Results
Table 3: Mean rank w.r.t. task-specific evaluation metric results on 18 ML tasks in the deployment stage. Results are reported over 10 repetitive runs. Best performances are highlighted in bold, and second best performances are underlined.
@@ 5 Related Work
LLM Agent. LLMs have demonstrated remarkable foundational capabilities, such as language understanding, complex reasoning, tool usage, and code generation, which gives rise to the development of autonomous language agents designed for various tasks (Yao et al., 2022; Hong et al., 2024; Wu et al., 2023; Wang et al., 2023c; Zhao et al., 2024; Boiko et al., 2023; Romera-Paredes et al., 2023; Deng et al., 2023; Lin et al., 2023). Within the field of data science, Mahadi Hassan et al. (2023) discusses the potential of LLMs as conversational agents in data science workflows. Moreover, recent studies have investigated the use of LLM agents in diverse areas such as feature engineering (Hollmann et al., 2023), hyper-parameter tuning (Zhang et al., 2023c, a), using ML libraries (Liu et al., 2023), aiding AI research (Huang et al., 2023), data operation (Lai et al., 2023), etc. In contrast to them, we focus on developing automatic language agents to build and train ML models, contributing to the field of automated data science.
@@ 6 Conclusion
Impact Statements
(This section does not count toward the paper page limit.)
@@
The appendix of this paper is organized as below. First, we provide the pseudo-code of DS-Agent in Appendix A. We then provide experimental details, including task selection (Appendix B.1), details for human insight collection (Appendix B.2), and model configuration and hyper-parameter settings (Appendix B.3). Subsequently, we provide further discussions on AutoML techniques (Appendix C.1). Moreover, we provide two case studies of DS-Agent in Appendix C.2. Finally, the detailed prompt design in DS-Agent is presented in Appendix D.
@@ B.1 Task Selection
Task Description
You are solving this machine learning tasks of regression:
The dataset presented here (the Airline reviews) comprises customer feedback for British Airways. Here, we provide the textual reviews. Your task is to predict the corresponding rating in the range of {1, ..., 10} given the reviews in the test set. The evaluation metric is root mean squared error (RMSE).
We provide an overall pipeline in train.py. Now fill in the provided train.py script to train a language model to get a good performance.
The provided Python script (train.py)
import pandas as pd from sklearn.metrics import mean_squared_error import numpy as np import random import torch from sklearn.model_selection import train_test_split from submission import submit_predictions_for_test_set SEED = 42 random.seed(SEED) torch.manual_seed(SEED) np.random.seed(SEED) device = torch.device("cuda" if torch.cuda.is_available() else "cpu") def compute_metrics_for_regression(y_test, y_test_pred):     rmse = mean_squared_error(y_test, y_test_pred, squared=False)     return rmse def train_model(X_train, y_train, X_valid, y_valid):     # TODO. define and train the model     # should return the trained model     model = None     return model def predict(model, X):     # TODO. predict the model     # should return an array of predictions     y_pred = np.random.randint(1, 11, len(X))     return y_pred if __name__ == ’__main__’:     data_df = pd.read_csv(’train.csv’)     data_df = data_df.dropna(subset=[’OverallRating’])     # Process data and store into numpy arrays.     X = list(data_df.ReviewBody.to_numpy())     y = data_df.OverallRating.to_numpy()     # Create a train-valid split of the data.     X_train, X_valid, y_train, y_valid = train_test_split(X, y, test_size=0.10, random_state=SEED)     # define and train the model     # should fill out the train_model function     model = train_model(X_train, y_train, X_valid, y_valid)     # evaluate the model on the valid set using compute_metrics_for_regression and print the results     # should fill out the predict function     y_valid_pred = predict(model, X_valid)     rmse = compute_metrics_for_regression(y_valid, y_valid_pred)     print("finalRMSEonvalidationset:", rmse)     # submit predictions for the test set     submission_df = pd.read_csv(’test.csv’)     submission_df = submission_df.dropna(subset=[’OverallRating’])     X_submission = list(submission_df.ReviewBody.to_numpy())     y_submission = predict(model, X_submission)     submit_predictions_for_test_set(y_submission)
@@ B.2 Details for Human Insight Case Collection
Prompt for solution extraction
Assume that you were a proficient data scientist. The following Python code is a high-performing solution for a kaggle competition.
Please answer the following questions one by one and **as detailedly as possible**.
Make sure that another data scientist can exactly reproduce this copy of code based on your answer.
Focus on the training process.
(1) Please give a summary of the overall design.
(2) What is the overall model architecture? Please use a long article to answer this question as accurately and in detail as possible.
(3) How are the important hyper-parameters setting in this code?
(4) What is the optimization objective?
(5) What advanced machine learning technique does this copy of code use?
(6) What other important tricks do you think play an important role for high performance?
Note that make sure the answers are directly included from the python code, rather than based on your assumption.
‘‘‘python
{Here is the Python code.}
An example human insight case that is derived from the public technical report
My solution is rather simple, because I used almost no complex tricks - instead I just built a reliable pipeline and found good hyperparameters. I had some brilliant ideas (at least I think so :)) - to use augmentation and generate synthetic data using some big model based on other texts from commonlit.org. But at that time, when I was still actively contributing, there was uncertainty with the license for the texts, and I did not complete the use of augmentations - I was too busy with my work (trying to get RLHF to work), so I left the competition - my last commit was a month ago. But apparently the decision was good, at least I didnt overfit :)
So lets break my solution down into parts.
**1. Data:**
I used a fairly standard template: prompt + question + text. At the end of my participation, I tried to dig into the data and do a good pre-processing - after all, there were quite a lot of very similar essays with exact or different scores. So, I tried to find these samples by similarity (like Levenstein) and then merge them. In addition, I decided to make augumentations based on this insight - if there are many similar essays (differing only in typos, for example) - I could use something like reverse autocorrect - randomly replace some words with its close analogues. With this technique I got 0.453 in private ONLY on fold3 (which is better than my chosen blend and probably could lead to a second place) - but I was too tired at this point so I didnt look further into augmentations. But I think augmentations could probably lead me to victory.
**2. Models**
Deberta is the king, so theres not much to say here. I tried using decoder models like Llama, but Deberta was still better. There were some techniques that gave me a boost - using EMA (honestly, without EMA it was very unstable, so its probably just necessary) and using differential learning rates.I tried several pooling options, but the best option for me was to use concatenation of CLS token and students text meanpooling. I also used token_type_ids to separate the prompt, question and essay.
**3 Inference & train**
I used following scheme - I tried to find a good hyperparameters on some fold (for example, fold0), and then train with exact hyperparameters on other folds. I then sumbittend the entire blend and 4 individual models (5 submission total - one day) and repeated the procedure the next day. I realized that I could use maxlen 1500 for inference (didnt research this number much, tried something like 1024 and 2048, but 1500 was better in terms of efficiency), so in my final mix I took the 10 best checkpoints across folds (some folds got 2 checkpoints, some folds got 3). First I averaged by folds, then averaged the rest. Thats all.
Briefly what worked (ranked from most important to least important, IMO):
1. Using Deberta
2. EMA
3. Augumentation
4. Defferentiated learning rates
5. Custom pooling
6. token_type_ids
7. Data cleaninig
What did not work (random order):
1. Decoder models
2. AWP
3. FGM
4. WD
5. Constant LR
6. Handcrafted features
7. GBT for stacking
In the end, it was a good competition for me. Last year I competed in another NLP competition and got a silver medal, but I grinded all day at that competition (I wasnt working that time, so I had a lot of free time). This time I also expected silver, which I consider a solid result, but I got 3rd place. In any case, this competition was a cakewalk for me, since I spend very little effort on it (compared to the previous competition, at least). Im hoping this means Ill grow a lot this year - and I think thats the main goal of participating in Kaggle.
Good luck to all of you.
An example human insight case that is derived from the public codes
(1) The overall design of the code is to train a DebertaV3 model for predicting the content and wording scores of student summaries. The code includes data preprocessing, model training, validation, and prediction steps.
(2) The overall model architecture is based on the DebertaV3 model, which is a transformer-based model. The code uses the AutoModelForSequenceClassification class from the transformers library to load the pre-trained DebertaV3 model. The model is fine-tuned for sequence classification with a single output label. The input to the model is a concatenation of the prompt question, summary text, and prompt text. The model tokenizes the input using the AutoTokenizer class and generates input tensors for the model. The model architecture consists of multiple transformer layers with self-attention mechanisms, followed by a linear layer for classification.
(3) The important hyperparameters in this code are set in the CFG class. The hyperparameters include the model name, learning rate, weight decay, hidden dropout probability, attention dropout probability, number of training epochs, number of cross-validation splits, batch size, random seed, save steps, and maximum sequence length.
(4) The optimization objective is to minimize the root mean squared error (RMSE) between the predicted scores and the ground truth scores. The code uses the mean squared error (MSE) as the loss function and calculates the RMSE as the evaluation metric.
(5) The advanced machine learning technique used in this code is transfer learning. The code loads a pre-trained DebertaV3 model and fine-tunes it on the student summary dataset. Transfer learning allows the model to leverage knowledge learned from a large pre-training dataset to improve performance on a specific task.
(6) Some important tricks that play a role in high performance include:
- Data preprocessing: The code preprocesses the input data by tokenizing the text, removing stop words, fixing misspellings, and extracting features such as text length, word overlap, n-gram co-occurrence, quotes overlap, and grammar check.
- Model architecture: The code uses the DebertaV3 model, which is a state-of-the-art transformer-based model known for its strong performance on various natural language processing tasks.
- Training strategy: The code uses k-fold cross-validation to train and validate the model on multiple subsets of the data. This helps to reduce overfitting and obtain a more robust evaluation of the models performance.
- Evaluation metric: The code uses the root mean squared error (RMSE) as the evaluation metric, which is a common metric for regression tasks. This metric penalizes large errors more than mean absolute error (MAE) and provides a more comprehensive measure of the models performance.
- Feature engineering: The code incorporates additional features such as word difficulty, readability scores, and cosine similarity between the summary and prompt text. These features capture different aspects of the text and can provide additional information for the model to make predictions.
- Ensemble learning: The code combines the predictions from multiple folds of the cross-validation to obtain a more robust prediction. This helps to reduce the variance and improve the overall performance of the model.

### v5

@@ Abstract
In this work, we investigate the potential of large language models (LLMs) based agents to automate data science tasks, with the goal of comprehending task requirements, then building and training the best-fit machine learning models. Despite their widespread success, existing LLM agents are hindered by generating unreasonable experiment plans within this scenario. To this end, we present DS-Agent, a novel automatic framework that harnesses LLM agent and case-based reasoning (CBR). In the development stage, DS-Agent follows the CBR framework to structure an automatic iteration pipeline, which can flexibly capitalize on the expert knowledge from Kaggle, and facilitate consistent performance improvement through the feedback mechanism. Moreover, DS-Agent implements a low-resource deployment stage with a simplified CBR paradigm to adapt past successful solutions from the development stage for direct code generation, significantly reducing the demand on foundational capabilities of LLMs. Empirically, DS-Agent with GPT-4 achieves 100% success rate in the development stage, while attaining 36% improvement on average one pass rate across alternative LLMs in the deployment stage. In both stages, DS-Agent achieves the best rank in performance, costing \$1.60 and \$0.13 per run with GPT-4, respectively. Our data and code are open-sourced at https://github.com/guosyjlu/DS-Agent.
@@ 1 Introduction
Empirically, we demonstrate the superiority of DS-Agent across 30 data science tasks in both stages. During the development stage, DS-Agent with GPT-4 achieves 100% success rate across 12 tasks. For the deployment stage, DS-Agent with GPT-3.5 and GPT-4 achieves 85% and 99% one pass rate over 18 deployment tasks, while the best baseline attains a mere 56% and 60%. Remarkably, DS-Agent improves the one pass rate of an open-source LLM Mixtral-8x7b-Instruct from a mere 6% to 31%. In both stages, DS-Agent with GPT-4 and GPT-3.5 attains the highest and second-highest ranks in performance. Moreover, DS-Agent costs \$0.06 and \$1.60 per run for GPT-3.5 and GPT-4 with standard scenario, which reduce further to \$0.0045 and \$0.135 in low-resource scenarios, rendering DS-Agent highly appealing for real-world deployment.
@@ 3.1 Development Stage: Automatic Iteration Pipeline
We summarize the pseudo-code of the automatic pipeline in Algorithm 1. Overall, DS-Agent benefits from the CBR paradigm in two aspects. Firstly, CBR integrates the human insight case bank, which contains intensive expert knowledge of the data science, enabling DS-Agent to derive reasonable experiment plans. Moreover, CBR offers a flexible learning mechanism by retaining successful solution cases into the human insight case bank, thus eliminating the need for resource-intensive finetuning of LLMs via back-propagation. For instance, when encountering novel tasks that involve previously unseen data modalities, such as graph data (Pei et al., 2020, 2024a, 2024b), it can simply integrate the latest human insights into the case bank $\mathcal{C}$. This enables DS-Agent to adeptly solve data science tasks related to graph data by drawing on its enriched knowledge repository.
@@ 4.1 Experiment Setting
Evaluation Metric. We mainly evaluate the agent’s ability from three aspects: (1) Completion of building ML models. In the development stage, we employ the success rate, i.e., whether the agent can build an ML model in a bug-free manner within a fixed number of steps. In the deployment stage, the metric is the one-pass rate, indicating the agent’s ability to build an ML model with only a single trial. (2) Performance of built ML models. For both stages, we utilize mean rank and best rank as the evaluation metric to evaluate the agents’ capabilities for automated data science. (3) Resource cost. Since we mainly utilize closed-source LLMs in this work, we take the consumed money to assess resource costs.
@@ 4.2.1 Main Results
Table 3: Mean rank w.r.t. task-specific evaluation metric results on 18 data science tasks in the deployment stage. Results are reported over 10 repetitive runs. Best performances are highlighted in bold, and second best performances are underlined.
@@ 5 Related Work
LLM Agent. LLMs have demonstrated remarkable foundational capabilities, such as language understanding, complex reasoning, tool usage, and code generation, which gives rise to the development of autonomous language agents designed for various tasks (Yao et al., 2022; Hong et al., 2024b; Wu et al., 2023; Wang et al., 2023c; Zhao et al., 2024; Boiko et al., 2023; Romera-Paredes et al., 2023; Deng et al., 2023; Lin et al., 2023). Within the field of data science, Mahadi Hassan et al. (2023) discusses the potential of LLMs as conversational agents in data science workflows. Moreover, recent studies have investigated the use of LLM agents in diverse areas such as feature engineering (Hollmann et al., 2023), hyper-parameter tuning (Zhang et al., 2023c, a), using ML libraries (Liu et al., 2023), aiding AI research (Huang et al., 2023), data operation (Lai et al., 2023), etc. In contrast to them, we focus on developing automatic language agents to build and train ML models, contributing to the field of automated data science. Concurrent to our work, Hong et al. (2024a) propose Data Interpreter, which focuses on optimizing the workflow of LLM agents for data science scenarios to fully unlock the intrinsic knowledge of LLMs. The core techniques of Data Interpreter and DS-Agent are complementary. A potential future work would be to enhance DS-Agent with Data Interpreter or integrate CBR into Data Interpreter for further improvement.
@@ 6 Conclusion
Acknowledgements
We truly thank the reviewers for their great effort in our submission. This work was supported by the National Key R&D Program of China under Grant (No. 2023YFF0905400), National Natural Science Foundation of China through grants (No. U2341229, No. 61976102, No. U19A2065), the Key R&D Project of Jilin Province, China, (No. 20240304200SF), and the International Cooperation Project of Jilin Province, China, (No. 20220402009GH).
Impact Statement
@@ Acknowledgements
The appendix of this paper is organized as below. First, we provide the pseudo-code of DS-Agent in Appendix A. We then provide experimental details, including task selection (Appendix B.1), details for human insight collection (Appendix B.2), and model configuration and hyper-parameter settings (Appendix B.3). Subsequently, we provide further discussions on AutoML techniques (Appendix C.1). Moreover, we provide two case studies of DS-Agent in Appendix C.2 and detailed error mode analyses in Appendix C.3. Finally, the detailed prompt design in DS-Agent is presented in Appendix D.
@@ B.1 Task Selection
Task Description You are solving this machine learning tasks of regression: The dataset presented here (the Airline reviews) comprises customer feedback for British Airways. Here, we provide the textual reviews. Your task is to predict the corresponding rating in the range of {1, ..., 10} given the reviews in the test set. The evaluation metric is root mean squared error (RMSE). We provide an overall pipeline in train.py. Now fill in the provided train.py script to train a language model to get a good performance.
The provided Python script (train.py) import pandas as pd from sklearn.metrics import mean_squared_error import numpy as np import random import torch from sklearn.model_selection import train_test_split from submission import submit_predictions_for_test_set SEED = 42 random.seed(SEED) torch.manual_seed(SEED) np.random.seed(SEED) device = torch.device("cuda" if torch.cuda.is_available() else "cpu") def compute_metrics_for_regression(y_test, y_test_pred): rmse = mean_squared_error(y_test, y_test_pred, squared=False) return rmse def train_model(X_train, y_train, X_valid, y_valid): # TODO. define and train the model # should return the trained model model = None return model def predict(model, X): # TODO. predict the model # should return an array of predictions y_pred = np.random.randint(1, 11, len(X)) return y_pred if __name__ == ’__main__’: data_df = pd.read_csv(’train.csv’) data_df = data_df.dropna(subset=[’OverallRating’]) # Process data and store into numpy arrays. X = list(data_df.ReviewBody.to_numpy()) y = data_df.OverallRating.to_numpy() # Create a train-valid split of the data. X_train, X_valid, y_train, y_valid = train_test_split(X, y, test_size=0.10, random_state=SEED) # define and train the model # should fill out the train_model function model = train_model(X_train, y_train, X_valid, y_valid) # evaluate the model on the valid set using compute_metrics_for_regression and print the results # should fill out the predict function y_valid_pred = predict(model, X_valid) rmse = compute_metrics_for_regression(y_valid, y_valid_pred) print("finalRMSEonvalidationset:", rmse) # submit predictions for the test set submission_df = pd.read_csv(’test.csv’) submission_df = submission_df.dropna(subset=[’OverallRating’]) X_submission = list(submission_df.ReviewBody.to_numpy()) y_submission = predict(model, X_submission) submit_predictions_for_test_set(y_submission)
@@ B.2 Details for Human Insight Case Collection
Prompt for solution extraction Assume that you were a proficient data scientist. The following Python code is a high-performing solution for a kaggle competition. Please answer the following questions one by one and **as detailedly as possible**. Make sure that another data scientist can exactly reproduce this copy of code based on your answer. Focus on the training process. (1) Please give a summary of the overall design. (2) What is the overall model architecture? Please use a long article to answer this question as accurately and in detail as possible. (3) How are the important hyper-parameters setting in this code? (4) What is the optimization objective? (5) What advanced machine learning technique does this copy of code use? (6) What other important tricks do you think play an important role for high performance? Note that make sure the answers are directly included from the python code, rather than based on your assumption. ‘‘‘python {Here is the Python code.} ‘‘‘
An example human insight case that is derived from the public technical report My solution is rather simple, because I used almost no complex tricks - instead I just built a reliable pipeline and found good hyperparameters. I had some brilliant ideas (at least I think so :)) - to use augmentation and generate synthetic data using some big model based on other texts from commonlit.org. But at that time, when I was still actively contributing, there was uncertainty with the license for the texts, and I did not complete the use of augmentations - I was too busy with my work (trying to get RLHF to work), so I left the competition - my last commit was a month ago. But apparently the decision was good, at least I didnt overfit :) So lets break my solution down into parts. **1. Data:** I used a fairly standard template: prompt + question + text. At the end of my participation, I tried to dig into the data and do a good pre-processing - after all, there were quite a lot of very similar essays with exact or different scores. So, I tried to find these samples by similarity (like Levenstein) and then merge them. In addition, I decided to make augumentations based on this insight - if there are many similar essays (differing only in typos, for example) - I could use something like reverse autocorrect - randomly replace some words with its close analogues. With this technique I got 0.453 in private ONLY on fold3 (which is better than my chosen blend and probably could lead to a second place) - but I was too tired at this point so I didnt look further into augmentations. But I think augmentations could probably lead me to victory. **2. Models** Deberta is the king, so theres not much to say here. I tried using decoder models like Llama, but Deberta was still better. There were some techniques that gave me a boost - using EMA (honestly, without EMA it was very unstable, so its probably just necessary) and using differential learning rates.I tried several pooling options, but the best option for me was to use concatenation of CLS token and students text meanpooling. I also used token_type_ids to separate the prompt, question and essay. **3 Inference & train** I used following scheme - I tried to find a good hyperparameters on some fold (for example, fold0), and then train with exact hyperparameters on other folds. I then sumbittend the entire blend and 4 individual models (5 submission total - one day) and repeated the procedure the next day. I realized that I could use maxlen 1500 for inference (didnt research this number much, tried something like 1024 and 2048, but 1500 was better in terms of efficiency), so in my final mix I took the 10 best checkpoints across folds (some folds got 2 checkpoints, some folds got 3). First I averaged by folds, then averaged the rest. Thats all. Briefly what worked (ranked from most important to least important, IMO): 1. Using Deberta 2. EMA 3. Augumentation 4. Defferentiated learning rates 5. Custom pooling 6. token_type_ids 7. Data cleaninig What did not work (random order): 1. Decoder models 2. AWP 3. FGM 4. WD 5. Constant LR 6. Handcrafted features 7. GBT for stacking In the end, it was a good competition for me. Last year I competed in another NLP competition and got a silver medal, but I grinded all day at that competition (I wasnt working that time, so I had a lot of free time). This time I also expected silver, which I consider a solid result, but I got 3rd place. In any case, this competition was a cakewalk for me, since I spend very little effort on it (compared to the previous competition, at least). Im hoping this means Ill grow a lot this year - and I think thats the main goal of participating in Kaggle. Good luck to all of you.
An example human insight case that is derived from the public codes (1) The overall design of the code is to train a DebertaV3 model for predicting the content and wording scores of student summaries. The code includes data preprocessing, model training, validation, and prediction steps. (2) The overall model architecture is based on the DebertaV3 model, which is a transformer-based model. The code uses the AutoModelForSequenceClassification class from the transformers library to load the pre-trained DebertaV3 model. The model is fine-tuned for sequence classification with a single output label. The input to the model is a concatenation of the prompt question, summary text, and prompt text. The model tokenizes the input using the AutoTokenizer class and generates input tensors for the model. The model architecture consists of multiple transformer layers with self-attention mechanisms, followed by a linear layer for classification. (3) The important hyperparameters in this code are set in the CFG class. The hyperparameters include the model name, learning rate, weight decay, hidden dropout probability, attention dropout probability, number of training epochs, number of cross-validation splits, batch size, random seed, save steps, and maximum sequence length. (4) The optimization objective is to minimize the root mean squared error (RMSE) between the predicted scores and the ground truth scores. The code uses the mean squared error (MSE) as the loss function and calculates the RMSE as the evaluation metric. (5) The advanced machine learning technique used in this code is transfer learning. The code loads a pre-trained DebertaV3 model and fine-tunes it on the student summary dataset. Transfer learning allows the model to leverage knowledge learned from a large pre-training dataset to improve performance on a specific task. (6) Some important tricks that play a role in high performance include: - Data preprocessing: The code preprocesses the input data by tokenizing the text, removing stop words, fixing misspellings, and extracting features such as text length, word overlap, n-gram co-occurrence, quotes overlap, and grammar check. - Model architecture: The code uses the DebertaV3 model, which is a state-of-the-art transformer-based model known for its strong performance on various natural language processing tasks. - Training strategy: The code uses k-fold cross-validation to train and validate the model on multiple subsets of the data. This helps to reduce overfitting and obtain a more robust evaluation of the models performance. - Evaluation metric: The code uses the root mean squared error (RMSE) as the evaluation metric, which is a common metric for regression tasks. This metric penalizes large errors more than mean absolute error (MAE) and provides a more comprehensive measure of the models performance. - Feature engineering: The code incorporates additional features such as word difficulty, readability scores, and cosine similarity between the summary and prompt text. These features capture different aspects of the text and can provide additional information for the model to make predictions. - Ensemble learning: The code combines the predictions from multiple folds of the cross-validation to obtain a more robust prediction. This helps to reduce the variance and improve the overall performance of the model.